# Smeared phase transition in the dissipative random quantum Ashkin-Teller model

### v1

@@ VII Discussion and conclusions
In this work, we investigated the effects of dissipation in the phase transitions of the quantum Ashkin-Teller chain in the presence of quenched disorder. The dissipationless model exhibits three quantum phase transitions (see Fig. 1(a)): a FM-PM transition, a FM-PROD transition, and a PM-PROD transtion. All belonging to the universality class of the random transverse-field Ising chain and displaying off-critical quantum Griffiths phases. Introducing dissipation of the quantum fluctuations, surprisingly, does not smear all of these phase transitions. Only the transitions involving the FM phase are smeared (see Fig. 1(b)).
@@
In the more familiar model of the random-transverse field Ising chain, the low-energy physics is dominated by the so-called rare regions in the Griffiths phases. These are regions which are locally FM and weakly coupled to the bulk. Their dynamics are, thus, slow coherent tunneling between the two lowest energy FM states. Smearing due to dissipation occurs because that slow dynamics is completely halted by Ohmic and sub-Ohmic dissipation. The coherent tunneling requires a coherent changing in all oscillators rest position. When that set of oscillators is sufficiently slow, the rare region stop tunneling and develops its own magnetic order regardless the bulk.
@@
It turns out that the slow tunneling of the rare regions in the PM-PROD phase does not require coherent tunneling of all their constituents. This is because the magnetic order of the PROD phase is a composite one: each color spin is disordered, $\left\langle S_{1}^{z}\right\rangle=\left\langle S_{2}^{z}\right\rangle=0$, while the composite magnetic order finite, $\left\langle S_{1}^{z}S_{2}^{z}\right\rangle\neq 0$. Thus, dissipation do not couple directly to the PROD order parameter and, therefore, the transition remains sharp.
@@
It is interesting to consider whether similar situations appear in other contexts. Composite (quadrupole-like) nematic order appears in diverse contexts such as in frustrated magnets (as, for instance, in the context of order-by-disorder in the $J_{1}$-$J_{2}$ Heisenberg model or in easy-plane pyrochlores) and in iron pnictides. Giving the results here uncovered, dissipation due to the coupling to gapless modes (such as that of a metallic Fermi liquid) may not couple to that composite magnetic order, and, consequently, the fluctuations of the associated nematic rare regions may not be overdamped by dissipation.
@@
It is also interesting to test our result in other models. In Appendix A we consider two possible modifications on the dissipation of the quantum fluctuations: one in which the couplings to the oscillator bath is color-dependent and another in which there is only one oscillator bath for both color spins. In both cases, our result remains: the smearing occurs only to the transitions to the FM phase.
@@
Finally, we argue that our results are not restricted to one spatial dimension. Our SDRG method can be applied to any dimension and the smearing (or lack of) will not depend on the underling coordination number.
@@ A.1 Color-dependent couplings
To check how dissipation will affect the phase transitions, we project the coupling Hamiltonian $H_{\text{SB}}$ onto the ground states of each decimation case. For cases (ii) and (iii), the coupling Hamiltonian have the same form as the color-independent case (apart from the color dependence of the $\lambda$’s). Therefore, there is no renormalization of the dissipation strength $\alpha$. For case (iv), proceeding analogously to Sec. VI, we get
@@ A.2 Single dissipation bath
Here, we investigate whether our results on the model Hamiltonian (5) also applies to another related dissipative model. In (5), each spin is coupled to its own set of oscillator bath. We now consider the case in which different color-spins are coupled to the same set of oscillators. Namely,
@@
where $x_{0;k,i}=\lambda_{k,i}/m_{k,i}\omega_{k,i}^{2}$ is the counter-term (caldeiraleggett). It must be added if we wish that the coupling to the bath introduces only dissipation, and not effective fields [which appears when completing the square in the last line of Eq. (34)]. To proceed with the adiabatic renormalization, we set
@@
and treat the tunneling terms of the $i$-th site in first order of perturbation theory. The definition of $\omega_{l}$ is the same as in Sec. IV. The 4-fold degenerated ground state of $H_{0}$ is given by
$\left|\Psi_{++}\right\rangle=$ $\left|+,+\right\rangle\prod_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;++}\right\rangle,$
$\left|\Psi_{-+}\right\rangle=$ $\left|-,+\right\rangle\prod_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;-+}\right\rangle,$ (36)
$\left|\Psi_{\pm-}\right\rangle=$ $\left|\pm,-\right\rangle\prod_{\omega_{k,i}>\omega_{l}}\left|0_{k,i}\right\rangle,$
where we are using the same notation as Sec. IV (first and indices to $\eta$ and $\sigma$ variables, respectively). The oscillators’ ground states are
@@
$-\tilde{g_{i}}\eta_{i}^{x}\left(\frac{1+\sigma_{i}^{z}}{2}\right)$ (39)
@@
From Eq. (39), it follows that the tunneling frequency $g_{i}$ splits in two parts: one that decreases under adiabatic renormalization and other that remains unaffected. This occurs because both states $\left|\Psi_{\pm-}\right\rangle$ are centered at the same position, such that the tunneling between states $\left|\Psi_{+-}\right\rangle$ and $\left|\Psi_{--}\right\rangle$, driven by the term $g_{i}\eta_{i}^{x}$ in $H_{1}$, does not change the equilibrium position of the oscillators. As a result, the tunneling frequency $g_{i}$ remains unaffected in this case. However, the $g_{i}\eta_{i}^{x}$ term also promotes tunneling between the states $\left|\Psi_{++}\right\rangle$ and $\left|\Psi_{-+}\right\rangle$, which are centered at different positions. For this tunneling process, the frequency $g_{i}$ undergoes a renormalization, as described in Eq. (41).

### v2

@@ VII Discussions and conclusions
In this work, we investigated the effects of dissipation in the phase transitions of the quantum Ashkin-Teller chain in the presence of quenched disorder. The dissipationless model exhibits three quantum phase transitions [see Fig. 1(a)]: a FM-PM transition, a FM-PROD transition, and a PM-PROD transition, all belonging to the universality class of the random transverse-field Ising chain and displaying off-critical quantum Griffiths phases. Introducing dissipation of the quantum fluctuations, surprisingly, does not smear all of these phase transitions. Only the transitions involving the FM phase are smeared [see Fig. 1(b)].
@@
In the more familiar model of the random transverse-field Ising chain, the low-energy physics is dominated by the so-called rare regions in the Griffiths phases. These are regions which are locally FM and weakly coupled to the bulk. Their dynamics are, thus, slow coherent tunneling between the two lowest energy FM states. Smearing due to dissipation occurs because that slow dynamics is completely halted by Ohmic and sub-Ohmic dissipation. The coherent tunneling requires a coherent changing in all oscillators’ rest position. When that set of oscillators is sufficiently slow, the rare region stops tunneling and develops its own magnetic order regardless of the bulk.
@@
It turns out that the slow tunneling of the rare regions in the PM-PROD phase does not require coherent tunneling of all their constituents. This is because the magnetic order of the PROD phase is a composite one: each color spin is disordered, $\left\langle S_{1}^{z}\right\rangle=\left\langle S_{2}^{z}\right\rangle=0$, while the composite magnetic order finite, $\left\langle S_{1}^{z}S_{2}^{z}\right\rangle\neq 0$. Thus dissipation does not couple directly to the PROD order parameter and, therefore, the transition remains sharp.
@@
It is interesting to consider whether similar situations appear in other contexts. Composite (quadrupolelike) nematic order appears in diverse contexts such as in frustrated magnets (as, for instance, in the context of order by disorder in the $J_{1}$-$J_{2}$ Heisenberg model or in easy plane pyrochlores) and in iron pnictides. Giving the results here uncovered that dissipation due to the coupling to gapless modes (such as that of a metallic Fermi liquid) may not couple to that composite magnetic order, and, consequently, the fluctuations of the associated nematic rare regions may not be overdamped by dissipation.
@@
It is also interesting to test our result in other models. In the Appendix A we consider two possible modifications on the dissipation of the quantum fluctuations: one in which the couplings to the oscillator bath are color dependent and another in which there is only one oscillator bath for both color spins. In both cases, our result remains: the smearing occurs only to the transitions to the FM phase.
@@
Finally, we argue that our results are not restricted to one spatial dimension. Our SDRG method can be applied to any dimension and the smearing (or lack thereof) will not depend on the underlying coordination number.
@@ A.1 Color-dependent couplings
To check how dissipation will affect the phase transitions, we project the coupling Hamiltonian $H_{\text{SB}}$ onto the ground states of each decimation case. For cases (ii) and (iii), the coupling Hamiltonian has the same form as the color-independent case (apart from the color dependence of the $\lambda$’s). Therefore, there is no renormalization of the dissipation strength $\alpha$. For case (iv), proceeding analogously to Sec. VI, we get
@@ A.2 Single dissipation bath
Here, we investigate whether our results on the model Hamiltonian (5) also apply to another related dissipative model. In (5), each spin is coupled to its own set of oscillator baths. We now consider the case in which different color-spins are coupled to the same set of oscillators. Namely,
@@
where $x_{0;k,i}=\lambda_{k,i}/m_{k,i}\omega_{k,i}^{2}$ is the counterterm [32]. It must be added if we wish that the coupling to the bath introduces only dissipation, and not effective fields [which appears when completing the square in the last line of Eq. (35)]. To proceed with the adiabatic renormalization, we set
@@
and treat the tunneling terms of the $i$th site in first order of perturbation theory. The definition of $\omega_{l}$ is the same as in Sec. IV. The fourfold degenerate ground state of $H_{0}$ is given by
$\left|\Psi_{++}\right\rangle=$ $\left|+,+\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;++}\right\rangle,$
$\left|\Psi_{-+}\right\rangle=$ $\left|-,+\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;-+}\right\rangle,$ (37)
$\left|\Psi_{\pm-}\right\rangle=$ $\left|\pm,-\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|0_{k,i}\right\rangle,$
where we are using the same notation as Sec. IV (first and second indices to $\eta$ and $\sigma$ variables, respectively). The oscillators’ ground states are
@@
$-\tilde{g_{i}}\eta_{i}^{x}\left(\frac{1+\sigma_{i}^{z}}{2}\right),$ (40)
@@
From Eq. (40), it follows that the tunneling frequency $g_{i}$ splits into two parts: one that decreases under adiabatic renormalization and another that remains unaffected. This occurs because both states $\left|\Psi_{\pm-}\right\rangle$ are centered at the same position, such that the tunneling between states $\left|\Psi_{+-}\right\rangle$ and $\left|\Psi_{--}\right\rangle$, driven by the term $g_{i}\eta_{i}^{x}$ in $H_{1}$, does not change the equilibrium position of the oscillators. As a result, the tunneling frequency $g_{i}$ remains unaffected in this case. However, the $g_{i}\eta_{i}^{x}$ term also promotes tunneling between the states $\left|\Psi_{++}\right\rangle$ and $\left|\Psi_{-+}\right\rangle$, which are centered at different positions. For this tunneling process, the frequency $g_{i}$ undergoes a renormalization, as described in Eq. (42).